OpenStudy (anonymous):

evaluate the function for the given value of x f(x)=4^x x=-2

OpenStudy (a_clan):

1/16

OpenStudy (anonymous):

f(x)=4^x x=-2 f(x)=4^(-2)=16^-1=1/16

OpenStudy (anonymous):

thank u...